# High-dimensional Laplace asymptotics up to the concentration threshold

### v1

@@ 1.2 Application in statistics
In this setting of optimizing an objective given by a high-dimensional integral, Laplace approximations are especially attractive. Indeed, they yield an explicit analytic surrogate for the objective as a function of the tuning parameter, rather than the black-box numerical access provided by methods such as Markov Chain Monte Carlo [33]. However, this convenience is only meaningful when the approximation error is controlled in a dimension-dependent way. In fact, concerns about the reliability of the Bayesian Information Criterion (BIC) in high dimensions have been noted explicitly in the literature. The BIC is a further approximation on top of the Laplace approximation used to simplify the optimization objective. In [10, p. 377], the authors write “for models with a large number of predictors [large $d$] it is no longer clear that some version of the BIC, or perhaps rather a Laplace approximation, accurately approximates a marginal likelihood”.
@@
While recent work has begun to provide rigorous guarantees for high-dimensional Laplace approximations [37, 3, 39, 22], these typically require more restrictive scalings (at best, that $d^{2}/\lambda\ll 1$). Our results push these guarantees essentially up to the concentration boundary, providing explicit “error bars” for Laplace-based surrogates deep into the modern high-dimensional regime.
@@ 1.3 Connection to cumulant theory
Our remainder bound makes rigorous the well-known observation that, based on the terms, expanding the cumulant generating function (cgf) is advantageous to expanding the moment generating function (mgf). Specifically, it is well-known that each cumulant is given by a sum of a fewer number of summands than the corresponding moment. For example at the end of Section 3.10.1 of [26], McCullagh writes that the formula for cumulants is a sum only over “connected pairs” of bi-partitions, as opposed to the formula for moments. In [16, Chapter 5.3], the author writes, “When calculating cumulants, only fully connected diagrams (without disjoint pieces) need to be included. This is a tremendous simplification.”
@@ 1.4 Related work
The paper is organized as follows. Section 2 introduces the notation and conventions used throughout. In Section 3, we describe the problem setting and state the main result, and Section 4 outlines the proof strategy. Section 5 derives the expansion of the main contribution to the integral coming from a neighborhood of zero, while Section 6 computes the expansion coefficients in terms of cumulants. The tail contribution is estimated in Section 7. Section 8 compares our results with the expansion of $I(\lambda)$ in [22] and presents two applications of our expansion. Various technical results are collected in the appendices.
@@ 3 Statement of result
    $|g(x)|e^{-\lambda f(x)}\leq c_{g}\exp\left(-\lambda\kappa\min\left\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|,d^{-1/(2L)}\log(1+\left\|x\right\|)\right\}\right),\quad\forall\,x\in\mathbb{R}^{d}.$ (3.6)
@@
In (3.6), recall that $\epsilon=d/\lambda$.
The case of a general minimizer $x_{\star}$ of $f$ and general positive definite $\nabla^{2}f(x_{\star})$ can be reduced to the one assumed above by an affine change of variables. By linearity, the normalization condition $g(0)=1$ is not restrictive. A sufficient condition for (3.6) to be satisfied is the following:
$\begin{split}f(x)&\geq\kappa\min\left\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|\right\},\quad\|x\|\leq r_{0},\\ f(x)&\geq\frac{1}{\lambda}\log|g(x)|+\kappa\min\left\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|,d^{-1/(2L)}\log(1+\left\|x\right\|)\right\},\quad\|x\|\geq r_{0}.\end{split}$ (3.7)
To see why (3.7) implies (3.6), note that the second line in (3.7) is equivalent to (3.6) in the domain $\|x\|\geq r_{0}$. Meanwhile, the first line is equivalent to $f(x)\geq\kappa\min\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|,d^{-1/(2L)}\log(1+\left\|x\right\|)\}$, $\|x\|\leq r_{0}$, for all $d^{L+1}/\lambda^{L}$ small enough. This can be combined with $\max_{\|x\|\leq r_{0}}|g(x)|\leq c_{g}$ to yield (3.6). The reason we can add the third argument to the minimum in the region $\|x\|\leq r_{0}$ is that $\sqrt{\epsilon}\|x\|\leq d^{-1/(2L)}\log(1+\left\|x\right\|)$ for all $\|x\|\leq r_{0}$ if $d^{L+1}/\lambda^{L}$ is small enough.
Consider the second, large-$\|x\|$ condition in (3.7). The logarithmic term in the minimum is the smallest of the three for large enough $x$. For such $x$ the condition becomes $f(x)\geq\frac{1}{\lambda}\log|g(x)|+\kappa d^{-1/(2L)}\log(1+\left\|x\right\|)$. Thus unless $g$ itself grows as some $g(x)=\exp(\|x\|^{p})$, $p>0$, we only require that $f$ grows logarithmically at infinity.
@@
The following is our main result.
@@
The coefficients $b_{k}(f,g)$ depend only on $\nabla^{\ell}f(0)$, $\ell=3,\dots,2L+1$, $\nabla^{\ell}\log g(0)$, $\ell=1,\dots,2L-1$, but not explicitly on $d$ or $\lambda$, and their formula is given in Lemma 4.3. Moreover, $|b_{k}(f,g)|\lesssim d^{k+1}$.
@@ 4.2 Roadmap
The coefficients $b_{k}(f,g,d)$ depend only on $d$, $\nabla^{\ell}f(0)$, $\ell=3,\dots,2L+1$, $\nabla^{\ell}\log g(0)$, $\ell=1,\dots,2L-1$, but not explicitly on $\lambda$. Moreover, $|b_{k}(f,g,d)|\lesssim d^{k+1}$.
@@
Here, $\mathrm{cum}(p_{\alpha}(Z))$ is defined at the beginning of Section 6.
Lemma 4.3 is proved in Section 6. Although we have used somewhat different notation, the above formula can be shown to coincide with that given in [26, Chapter 3.10.2].
@@
Lemma 4.5 is proved in Section 7. The three lemmas conclude the proof, since they give the desired form for the coefficients, and we have
@@
The integrand must be positive for (5.2) to hold. But indeed, $g(0)=1$ and $\log g$ has bounded first derivative, both by ((2)) of Assumption 3.1. Thus we can ensure $g(x)>0$ for all $\|x\|\leq R\sqrt{\epsilon}$ by choosing the upper bound $c^{+}$ on $R\sqrt{\epsilon}$ small enough in (4.9).
@@ 5.1 Iterative change of variables
Let $f_{2L+1}$ be as in (3.2). There exists $\varphi(t)=\sum_{q\geq 2}F_{q\to 1}^{\epsilon}[t^{\otimes q}]$ such that $f_{2L+1}(t+\varphi(t))=\frac{1}{2}\|t\|^{2}+\mathcal{O}(\|t\|^{2L+2})$ for all $\|t\|\leq 2R\sqrt{\epsilon}$.
See Appendix B for the proof. We now show that $\mathrm{id}+\varphi$ is bijective, and characterize the set $(\mathrm{id}+\varphi)^{-1}(\{\|x\|\leq R\sqrt{\epsilon}\})$. The following lemma states a slightly more general result which will be needed later on.
@@
Let $C$ be an absolute constant to be chosen later. There exists a change of variables
$T_{m}(s)=s+\epsilon^{m}\varphi_{m}(s),\quad\varphi_{m}(s)=\sum_{k\geq 2}F_{k\to 1}^{\epsilon}[s^{\otimes k}],$ (5.12)
@@
See the end of Appendix B for the proof. The value of $C$ will be chosen below the proof of Corollary 5.9.
@@
Thus we see that in Lemma5.7, we can take $C=2^{L}$. Note that (5.21) holds also for $L=1$.
@@ 5.2 Completing the square
$\log I_{\mathrm{in}}(\lambda)=\frac{1}{2}\lambda\epsilon^{2}{F_{0\to 1}^{\epsilon}}^{\top}(I_{d}+2\epsilon F_{0\to 2}^{\epsilon})^{-1}F_{0\to 1}^{\epsilon}-\frac{1}{2}\log\det(I_{d}+2\epsilon F_{0\to 2}^{\epsilon})+\mathcal{O}(R^{2L+2}d\epsilon^{L}).$ (5.22)
for two composite tensors $F_{0\to 1}^{\epsilon},F_{0\to 2}^{\epsilon}$.
Remark 5.11.
The constant vector $F_{0\to 1}^{\epsilon}$ is the same as the vector representation of $F_{1}^{\epsilon}$ in (5.20). The constant matrix $F_{0\to 2}^{\epsilon}$ is the same as the matrix representation of $F_{2}^{\epsilon}$ in (5.20). However, this fact is immaterial since we have not been tracking any explicit formulas. We only need to know the structure of the tensors.
@@
When $L\geq 2$, the function $E_{L}(t)$ is as in (5.20), containing a linear and quadratic correction to $\|t\|^{2}/2$. When $L=1$, then $E_{1}(t)$ from (5.10) has only a linear correction. Thus both cases can be considered to be of the form of $E_{L}$ from (5.20), since zero is a special case of a quadratic correction. Let us identify $F_{1}^{\epsilon}$ and $F_{2}^{\epsilon}$ in the definition of $E_{L}$ with $F_{0\to 1}^{\epsilon}$ and $F_{0\to 2}^{\epsilon}$ in the natural way, respectively. Since $F_{2}^{\epsilon}$ is a symmetric linear form, the corresponding $F_{0\to 2}^{\epsilon}$ is a symmetric matrix. Define the vector $a$ and matrix $B$ as
$a=\epsilon F_{0\to 1}^{\epsilon},\quad B=I_{d}+2\epsilon F_{0\to 2}^{\epsilon},$ (5.23)
and note that $B$ is also symmetric. Then
@@
The matrix $B$ is invertible if $\epsilon$ is small enough. We now do a final change of variables. Let $s=B^{1/2}t+B^{-1/2}a$. Then
@@
To conclude, substitute the definition of $a$ and $B$ from (5.23). ∎
@@ 5.3 Structure of terms in the expansion
Consider the two terms in (5.22). For brevity, we write $J_{1}=F_{0\to 1}^{\epsilon}$ and $J_{2}=F_{0\to 2}^{\epsilon}$ for the two specific tensors arising in (5.22). We have $\log\det(I_{d}+2\epsilon J_{2})=\mathrm{tr}\log(I_{d}+2\epsilon J_{2})$. Using that $\|2\epsilon J_{2}\|_{\mathrm{op}}<1/2$ if $\epsilon$ is sufficiently small (since $J_{2}=F_{0\to 2}^{\epsilon}$), we have
@@
From the Definition 5.1 and since $J_{1}=F_{0\to 1}^{\epsilon}$, it is clear that $J_{1}^{\top}F_{0\to 2}^{\epsilon}J_{1}=F_{0\to 0}^{\epsilon}=\mu$. In other words, $\mu$ is just a polynomial in $\epsilon$, i.e. $\mu=\sum_{k\geq 0}\mu_{k}\epsilon^{k}$. Its coefficients are bounded and depend only on $d$, $\nabla^{k}f(0)$, and $\nabla^{k}(\log g)(0)$. We conclude that

### v2

@@ 1.2.1 Statistics context and state of the art
In contrast, LE-based approximations are widely used for model selection problems that maximize the normalizing constant (model evidence) $\int e^{-\lambda f(x)}\mathrm{d}x$ over a tuning parameter. A key attraction is that the LE provides an explicit analytic surrogate objective as a function of the tuning parameter, enabling efficient optimization. This stands in sharp contrast to methods that typically provide only black-box numerical access to the objective, such as Markov Chain Monte Carlo [42]. However, this convenience of the LE is meaningful only when the approximation error is controlled in a dimension-dependent manner. In particular, concerns about the reliability of the Bayesian Information Criterion (BIC) in high dimensions have been noted explicitly. The BIC introduces an additional approximation on top of the LE to further simplify the optimization objective. In [13, p. 377], the authors write “for models with a large number of predictors [large $d$] it is no longer clear that some version of the BIC, or perhaps rather a Laplace approximation, accurately approximates a marginal likelihood”.
@@
While recent work has begun to provide rigorous, dimension-dependent guarantees for high-dimensional LEs for normalizing constants [3, 46, 28], these require more restrictive scalings (at best, that $d^{2}/\lambda\ll 1$).
1.2.2 Implication of our results in statistics
Our work essentially completes the program of finding tractable analytic approximations of posterior densities, expectations, and normalizing constants in the high-dimensional but concentrating regime. We propose a combined approach for these related problems which allows dimension to grow relative to $\lambda$ arbitrarily close to the concentration threshold.
For sampling, we construct a sequence of arbitrarily accurate approximations $\hat{\pi}_{L}$ which can be sampled from using an explicit algorithm. For computing expectations of smooth functions, we give an arbitrarily accurate closed-form formula based on the LE of numerator and denominator. For normalizing constants, (1.2) directly applies with $g\equiv 1$. The approximation error for all these tasks is $d^{L+1}/\lambda^{L}$, which has the crucial feature that increasing $L$ not only improves the accuracy but also expands the range of applicable $d$, since we can take $d$ as large as $o(\lambda^{L/L+1})$.
We now give more detail about our methods.
Sampling and nonsmooth expectations.
To approximately sample from $\pi$, the algorithm consists of drawing $Z_{i}$ from $\mathcal{N}(0,\lambda^{-1}I_{d})$ and mapping it through one of the $x_{L}$, $L=1,2,3,\dots$, depending on the desired accuracy. To approximate expectations of nonsmooth functions $g$, the samples $X_{i}=x_{L}(Z_{i})\sim\hat{\pi}_{L}$ can also be used, via
$\mathbb{E}_{X\sim\pi}[g(X)]\approx\mathbb{E}_{X\sim\hat{\pi}_{L}}[g(X)]\approx\frac{1}{N}\sum_{i=1}^{N}g(X_{i}).$ (1.11)
Normalizing constants and smooth expectations.
Approximating normalizing constants is immediate using (1.2) with $g\equiv 1$, so we don’t discuss it further. To approximate expectations $\mathbb{E}_{X\sim\pi}[g(X)]=\int ge^{-\lambda f}/\int e^{-\lambda f}$ of smooth $g$, we use (1.2) for the numerator and denominator. This gives the approximation $\mathbb{E}_{X\sim\pi}[g(X)]\approx\exp\bigl(\sum_{k=1}^{L-1}[b_{k}(f,g)-b_{k}(f,1)]/\lambda^{k}\bigr)$, which has accuracy $\mathcal{O}(d^{L+1}/\lambda^{L})$ uniformly over sufficiently smooth $g$ with bounded derivatives near $x_{\star}$.
This estimate improves on (1.11) in two ways. First, it does not incur any sampling error. Second, it is constructed using fewer derivatives of $f$: 2L12L-1 versus $2L+1$ for $\hat{\pi}_{L}$. Intuitively, $\hat{\pi}_{L}$ does not exploit smoothness of $g$ and must compensate with more information from $f$. In practice, this matters because $f$ encodes the data likelihood and is expensive to differentiate, while the observable $g$ is typically simple. For example, when $L=2$, the closed-form estimate $\exp([b_{1}(f,g)-b_{1}(f,1)]/\lambda)$ has accuracy $d^{3}/\lambda^{2}$ using only three derivatives of $f$, whereas $\hat{\pi}_{1}$ also uses three derivatives but achieves only the lower accuracy $d^{2}/\lambda$. Thus using the closed-form asymptotic estimate improves the accuracy by a factor of $d/\lambda$ over sampling approaches without requiring any more derivatives. We emphasize that most approaches in the literature are of the sampling type and achieve accuracy $d^{2}/\lambda$ (our $\hat{\pi}_{1}$ and [25]) or $d^{3}/\lambda$ [14] at best, even disregarding the Monte Carlo error. See Section 9.4 for a detailed comparison to the two works closest to ours, [14, 25].
@@ 1.3 Connection to cumulant theory
Our remainder bound makes rigorous the well-known observation that, based on the terms, expanding the cumulant generating function (cgf) is advantageous to expanding the moment generating function (mgf). Specifically, it is well-known that each cumulant is given by a sum of a fewer number of summands than the corresponding moment. For example at the end of Section 3.10.1 of [34], McCullagh writes that the formula for cumulants is a sum only over “connected pairs” of bi-partitions, as opposed to the formula for moments. In [21, Chapter 5.3], the author writes, “When calculating cumulants, only fully connected diagrams (without disjoint pieces) need to be included. This is a tremendous simplification.”
@@ 1.4 Related work
The paper is organized as follows. Section 2 introduces the notation and conventions used throughout. In Section 3, we describe the problem setting and state the main result, and Section 4 outlines the proof strategy. Section 5 derives the expansion of the main contribution to the integral coming from a neighborhood of zero, while Section 6 computes the expansion coefficients in terms of cumulants. The tail contribution is estimated in Section 7. Section 8 presents our results on the approximation of $\pi\propto e^{-\lambda f}$ and on computing expectations against $\pi$. Section 9 compares our results with related work in the literature, and presents two applications of our expansion. Various technical results are collected in the appendices.
@@ 3 Statement of result
    $|g(x)|e^{-\lambda f(x)}\leq\exp\left(c_{g}-\lambda\kappa\min\left\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|,d^{-1/(2L)}\log(1+\left\|x\right\|)\right\}\right),\quad\forall\,x\in\mathbb{R}^{d}.$ (3.6)
@@
In (3.6) and everywhere below, recall that $\epsilon=d/\lambda$.
The case of a general minimizer $x_{\star}$ of $f$ and general positive definite $\nabla^{2}f(x_{\star})$ can be reduced to the one assumed above by an affine change of variables. By linearity, the normalization condition $g(0)=1$ is not restrictive. A sufficient condition for (3.6) to be satisfied is
$f(x)$ $\geq\kappa_{f}\min\left\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|,d^{-1/(2L)}\log(1+\left\|x\right\|)\right\},\quad x\in\mathbb{R}^{d},$ (3.7)
$|g(x)|$ $\leq\min\left\{e^{\kappa_{g}\lambda\|x\|^{2}/2},e^{\kappa_{g}\sqrt{d\lambda}\|x\|},(1+\|x\|)^{\lambda\kappa_{g}/d^{1/(2L)}}\right\},\quad\|x\|\geq r_{0},$ (3.8)
$\kappa$ $:=\kappa_{f}-\kappa_{g}>0.$ (3.9)
When $x$ is small, the minimum in the first line is equal to $\min\{\|x\|^{2}/2,\sqrt{\epsilon}\left\|x\right\|\}$. When $x$ is large, the minimum is given by $d^{-1/(2L)}\log(1+\left\|x\right\|)$. Thus we only require that $f$ grows logarithmically at infinity. Meanwhile, $g$ can grow at most polynomially at infinity under (3.8), but the power of the polynomial can become arbitrarily large as $\lambda\to\infty$.
@@
The following is our first main result.
@@
The coefficients $b_{k}(f,g)$ depend only on $\nabla^{\ell}f(0)$, $\ell=3,\dots,2k+2$, $\nabla^{\ell}\log g(0)$, $\ell=1,\dots,2k$, but not explicitly on $d$ or $\lambda$, and their formula is given in Lemma 4.3. Moreover, $|b_{k}(f,g)|\lesssim d^{k+1}$.
@@ 4.2 Roadmap
The coefficients $b_{k}(f,g,d)$ do not explicitly depend on $\lambda$ and satisfy $|b_{k}(f,g,d)|\lesssim d^{k+1}$.
@@
Here, $\mathrm{cum}(p_{\alpha}(Z))$ is defined at the beginning of Section 6. Furthermore, $b_{M/2}(f,g)$ depends on derivatives of $f$ of order $3,\dots,M+2$ and derivatives of $\log g$ of order $1,\dots,M$.
Lemma 4.3 is proved in Section 6. Although we have used somewhat different notation, the above formula can be shown to coincide with that given in [34, Chapter 3.10.2].
@@
Lemma 4.5 is proved in Section 7. The three lemmas conclude the proof of Theorem 3.2. Indeed, they imply $I(\lambda)>0$ when $R^{2L}d^{L+1}/\lambda^{L}$ is sufficiently small, give the desired form for the coefficients, and imply
@@
The integrand must be positive for (5.2) to hold. But indeed, $g(0)=1$ and $\log g$ has bounded first derivative, both by (2) of Assumption 3.1. Thus we can ensure $g(x)>0$ for all $\|x\|\leq R\sqrt{\epsilon}$ by choosing the upper bound $c^{+}$ on $R\sqrt{\epsilon}$ small enough in (4.9). This proves our claim in Lemma 4.2 that $I_{\mathrm{in}}(\lambda)>0$.
@@ 5.1 Iterative change of variables
Let $f_{2L+1}$ be as in (3.2). There exists $X(t)=t+\varphi(t)$ with $\varphi(t)=\sum_{q=2}^{2L}F_{q\to 1}^{\epsilon}[t^{\otimes q}]$ such that $f_{2L+1}(X(t))=\frac{1}{2}\|t\|^{2}+\mathcal{O}(\|t\|^{2L+2})$ for all $\|t\|\leq 2R\sqrt{\epsilon}$. The function $\varphi$ is explicitly computable from the derivatives of $f$ of order $3,\dots,2L+1$.
See Appendix B for the proof and construction of $\varphi$. As an example, when $L=1$, we take $X(t)=t-\frac{1}{6}\nabla^{3}f(0)[t^{\otimes 2}]$, where $\nabla^{3}f(0)[t^{\otimes 2}]$ is the vector such that $\nabla^{3}f(0)[t^{\otimes 2}]^{\top}u=\nabla^{3}f(0)[t,t,u]$ for all $u\in\mathbb{R}^{d}$. It is then straightforward to check that
$\begin{split}f_{3}\left(t-\tfrac{1}{6}\nabla^{3}f(0)[t^{\otimes 2}]\right)&=\tfrac{1}{2}\left\|t-\tfrac{1}{6}\nabla^{3}f(0)[t^{\otimes 2}]\right\|^{2}+\tfrac{1}{6}\nabla^{3}f(0)\left[\left(t-\tfrac{1}{6}\nabla^{3}f(0)[t^{\otimes 2}]\right)^{\otimes 3}\right]\\ &=\|t\|^{2}/2+\mathcal{O}(\|t\|^{4}).\end{split}$ (5.3)
The basic observation that substituting $x=t-F[t^{\otimes k-1}]$ into $\|x\|^{2}/2+F[x^{\otimes k}]$ kills the order $k$ polynomial is at the heart of all of our changes of variables.
We now show that $X$ is bijective and characterize the set $X^{-1}(\{\|x\|\leq R\sqrt{\epsilon}\})$. The following lemma states a slightly more general result which will be needed later on.
@@
Let $C$ be an absolute constant to be chosen later. There is an explicitly computable change of variables
$T_{m}(s)=s+\epsilon^{m}\varphi_{m}(s),\quad\varphi_{m}(s)=\sum_{k=2}^{2L-2m}F_{k\to 1}^{\epsilon}[s^{\otimes k}],$ (5.13)
@@
See the end of Appendix B for the proof and the construction of $T_{m}$. The value of $C$ will be chosen below the proof of Corollary 5.9.
@@
Thus we see that in Lemma 5.7, we can take $C=2^{L}$. Note that (5.21) and (5.22) both also hold for $L=1$, i.e. $E_{L}$ from (5.21) coincides in structure with $E_{1}$ from (5.11), upon setting $F_{2}^{\epsilon}=0$ in (5.21).
The following quantities will appear in Sections 5.25.3, and 8.
Definition 5.10.
Let $a,B$ be such that $E_{L}$ in (5.21) can be written as
$E_{L}(t)=a^{\top}t+\frac{1}{2}t^{\top}Bt.$ (5.23)
Thus in particular,
$\begin{split}a&=\epsilon J_{1},\quad B=I_{d}+2\epsilon J_{2},\\ J_{1}&=F_{0\to 1}^{\epsilon},\quad J_{2}=F_{0\to 2}^{\epsilon}.\end{split}$ (5.24)
where $F_{0\to 1}^{\epsilon}$ and $F_{0\to 2}^{\epsilon}$ are the vector and matrix identifications of the specific $F_{1}^{\epsilon}$, $F_{2}^{\epsilon}$ appearing in $E_{L}$ in (5.21), respectively.
Since $F_{2}^{\epsilon}$ is by definition a symmetric bilinear form, the matrix $B$ is symmetric.
@@ 5.2 Completing the square
$\log I_{\mathrm{in}}(\lambda)=\frac{\lambda}{2}a^{\top}B^{-1}a-\frac{1}{2}\log\det B+\mathcal{O}(R^{2L+2}d\epsilon^{L}),$ (5.25)
where $a,B$ are as in (5.24).
@@
Recall $B$ is symmetric, and it is invertible if $\epsilon$ is small enough. By definition of $a,B$, we have
@@
We now do the final change of variables
$t=Q(s)=B^{-1/2}(s-B^{-1/2}a).$ (5.27)
Then
@@
This proves (5.25).∎
@@ 5.3 Structure of terms in the expansion
Recall from (5.24) that $a=\epsilon J_{1}$ and $B=I_{d}+2\epsilon J_{2}$, and consider the two terms in (5.25). We have $\log\det(I_{d}+2\epsilon J_{2})=\mathrm{tr}\log(I_{d}+2\epsilon J_{2})$. Using that $\|2\epsilon J_{2}\|_{\mathrm{op}}<1/2$ if $\epsilon$ is sufficiently small (since $J_{2}=F_{0\to 2}^{\epsilon}$), we have
@@
From the Definition 5.1 and since $J_{1}=F_{0\to 1}^{\epsilon}$, it is clear that $J_{1}^{\top}F_{0\to 2}^{\epsilon}J_{1}=F_{0\to 0}^{\epsilon}=:\mu$, where $\mu$ is just a polynomial in $\epsilon$, i.e. $\mu=\sum_{k\geq 0}\mu_{k}\epsilon^{k}$. Its coefficients are bounded and depend only on $d$, $\nabla^{k}f(0)$, and $\nabla^{k}(\log g)(0)$. We conclude that